Solve the initial value problem below using the method of Laplace transforms.

y''+8y'+15y=594e^(6t)

,y(0)=-4,y'(0)=78
Answer #1

Cliven Heit : y+8y+ 154 = 599.66t, 700)=-4 -0 y 10-78 het h be the Laplace oferecta. then fly) = $?.763) – Sy(0) – y()2463)885) (883) = 594- 482-243-468 +276 (56) YE). 870-482-225 (3-6) (315) (:S4.3) Mow use : penitical practia 070-44-228 (8-6sohe 5A - 6.03-10 c = 290- 670 At Beca.q. I . 8A-3B-C.=-22. 19A-23 = -261 AFB+C = -4 x(+10) 1. SA-6B-106 = 290 x 6 TO 10 X +1soby ④ x 99 -.B=-26 16232-26 – 901 . ?. 76 1.2) 2B = 26 + 901 .. 5 1167901 = 1312 4116 +901 - 1212 B: 1317 32. Afece 198 +131put in yes (5-6) (545) (543) Hlow telle : Inverse laplace transform * (YB)) = 6616) tui ti (hts) - si ti () 3517(4) = 6.26